OpenStudy (anonymous):

i ham having trouble finding the common denominator for 17 & 12

OpenStudy (anonymous):

there is no common denominator for 17 & 12

jhonyy9 (jhonyy9):

4x17=68x3=204

OpenStudy (anonymous):

check out the nos,its 17 & 12?

jhonyy9 (jhonyy9):

will be 204

OpenStudy (amistre64):

\[\frac{1}{17}+\frac{1}{12}=D\] \[\frac{1(12)}{17}+\frac{1(12)}{12}=(12)D\] \[\frac{12}{17}+1=12D\] \[\frac{12(17)}{17}+1(17)=12(17)D\] \[12+17=204D\] common denom is 204